# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

### v5

@@ Appendix A.2 Extended results on NExT-QA and IntentQA
We present extended zero-shot evaluation results on NExT-QA in Table A.7, comparing LVNet with prior zero-shot models across different task categories: causal, temporal, and descriptive reasoning. Models are ordered based on the number of captions processed per video, highlighting the trade-offs between caption efficiency and performance.
@@
We present extended zero-shot evaluation results on IntentQA in Table A.8, comparing LVNet with prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Before/After). Models are ordered based on the number of captions processed per video, highlighting the balance between caption efficiency and performance.
@@
To ensure fairness, models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in grey or downplayed in light green in Table A.7 and Table A.8. We adopt the reported GPT-4o results for VideoAgent and VideoTree in Section 4 and LABEL:tab:ablation:numberofframecaptions from the VCA Yang et al. (2024), but do not compare directly against VCA for two reasons: (1) the VCA paper does not provide code or implementation details (e.g., inference speed), making replication infeasible; and (2) its reported results cover only a subset of Egoschema, preventing a fair comparison to our approach on the full-scale EgoSchema, NExT-QA, and IntentQA. Overeall, these clarifications, alongside the results in Table A.7 and Table A.8, underscore LVNet’s effectiveness in achieving high accuracy while maintaining computational efficiency.
@@
SeViLA (Yu et al., 2024) 32 61.3 61.5 75.6 63.6
VideoTree (Wang et al., 2024e) (56) 75.2 67.0 81.3 73.5
@@
SeViLA (Yu et al., 2024) 32 - - - 60.9
VideoTree (Wang et al., 2024e) (56) - - - 66.9
@@
Table A.8: Extended results on IntentQA (Li et al., 2023a). We compare LVNet against prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Belief/Action). LVNet achieves an overall accuracy of 71.7%, surpassing all models while using only 12 captions per video. It reaches 75.0% in the Why? category, 74.4% in the How? category, and 62.1% in the B.A. category. Compared to VideoTree, which processes 56 captions and achieves 66.9% accuracy, LVNet outperforms it by +4.8% while using significantly fewer captions. Additionally, LVNet demonstrates superior reasoning-based performance compared to LangRepo (90 captions, 59.1%) and LLoVi (90 captions, 64.0%). Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency. Numbers in parentheses () indicate the maximum number of frames used.
@@ Appendix A.3 Algorithms in Detail
Our algorithms are presented in full detail in Algorithm 1, Algorithm 2, and Algorithm 3. TSC in Algorithm 1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the $n$ sets, we uniformly sample $\leq\tau$ frames, obtaining a total of $T_{a}\leq\tau\times n$ frames. For example, LVNet sets $\psi=5,\lambda=12,\tau=18$, resulting in approximately $n\sim 25$ and $T_{a}\sim 390$ on the EgoSchema dataset. CKD in Algorithm 2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16. LVNet employs $L=32,len(K)\leq 25$ on the EgoSchema dataset. FKD in Algorithm 3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally. It groups frames sequentially into visual templates, each consisting of $N$ frames. From each template, the $M$ frames and keywords most relevant among the $N$ pairs are selected using GPT-4o. We set $L=32,K=16,N=8,M=3$.
@@
43:   $\mathbf{F}^{i}\leftarrow f(img^{i})$
4:   $\textbf{List}_{feat}$.insert($\mathbf{F}^{i}$)
@@
67:   $\textbf{List}_{dist}\leftarrow\frac{\sum_{y}\sum_{x}{\sqrt{(\mathbf{F_{i}}-\textbf{List}_{feat})^{2}}}}{x\times y}$
8:   $\textbf{M}_{dist}$.insert($\textbf{List}_{dist}$)
@@
11:   $\textbf{List}_{sample}\leftarrow\emptyset$
12:   $\textbf{List}_{\delta}\leftarrow\emptyset$
813:   $i\leftarrow\textbf{List}_{index}$.pop(00)
914:   $\textbf{p}^{i}\leftarrow\texttt{softmax}(\textbf{M}^{i}_{dist})$
1015:   $\mu_{\mathbf{\textbf{p}^{i}}},\sigma_{\mathbf{\textbf{p}^{i}}}\leftarrow\texttt{mean}(\textbf{p}^{i}),\texttt{std}(\textbf{p}^{i})$
1116:   $\beta\leftarrow\mu_{\mathbf{\textbf{p}^{i}}}-\sigma_{\mathbf{\textbf{p}^{i}}}\sum_{i=0}{e^{1-i/\lambda}}$
@@
19:         $\textbf{List}_{selected}$.insert($\textbf{index}(\texttt{prob},\textbf{p}^{i})$)
@@
1323:      $\delta$leftarrow $\gamma~{}th$ value in $\textbf{List}_{index}$
@@
1527:   $\textbf{List}_{\delta}$.insert($i$)
1628:   $\textbf{List}_{sample}$leftarrow sample $\tau$ items from $\textbf{List}_{\delta}$
1729:   $\textbf{sort}(\textbf{List}_{sample})$
@@
42:$\mathbf{S}\leftarrow\textbf{sim}(\mathbf{K},\mathbf{I})$
53:$\mathbf{S}_{sorted},\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
@@
3
42:$\mathbf{idx}_{sorted}\leftarrow\textbf{sort}(\mathbf{S})$
@@
64:$\mathbf{idx}_{temporal}\leftarrow\textbf{temporal\_order}(\mathbf{idx}_{low\_sim})$
@@
86:$\mathbf{I}_{ordered},\mathbf{K}_{ordered}\leftarrow\mathbf{I}[\mathbf{idx}_{final}],\mathbf{K}[\mathbf{idx}_{final}]$
@@
9:   $\mathbf{I}_{selected}\leftarrow\texttt{select\_top\_M}(\mathbf{set},M)$
@@
$1$:return $\mathbf{I}_{selected}$
@@ Appendix A.4 Prompting: Fine Keyframe Detector
We prompt the VLM to select frames that are most compatible with the list of given keywords. Each template image contains 8 images, and their order is described in language (e.g. top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure A.7.
@@ Appendix A.5 Comparison with Other Keyframe Selection Methods
Figure A.5 visualizes the differences of the keyframe selection mechanism bewtween LVNet and VideoAgent. On the left, LVNet begins with uniformly sampled frames and filters them through multiple stages, resulting in a non-uniform distribution of frames over time. First, the temporal scene clustering (TSC) selects somes frames that represent temporally distinct activities. Next, the coarse keyframe detector (CKD) targets frames most relevant to the question. Finally, the fine keyframe detector (FKD) further refines this selection to ensure the keyframes accurately capture the activity in question. As a result, LVNet produces 12 frames, with 8 of them (67%) directly depicting "usage of phones," which is the correct answer and leads the model to select the right option. On the right, VideoAgent also starts with the uniform frames but relies on a LLM to request additional frames. Since the initial frames do not capture enough relevant content, the LLM again selects frames uniformly, adding more irrelevant samples that lack the crucial information about "usage of phones." As a result, VideoAgent ultimately selects the wrong option.

### v6

@@ Appendix A.2 Extended results on NExT-QA and IntentQA
We present extended zero-shot evaluation results on NExT-QA in Table˜A.12, comparing LVNet with prior zero-shot models across different task categories: causal, temporal, and descriptive reasoning. Models are ordered based on the number of captions processed per video, highlighting the trade-offs between caption efficiency and performance.
@@
We present extended zero-shot evaluation results on IntentQA in Table˜A.13, comparing LVNet with prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Before/After). Models are ordered based on the number of captions processed per video, highlighting the balance between caption efficiency and performance.
@@
To ensure fairness, models that utilize video-caption pretraining or process substantially more captions than LVNet are de-emphasized in grey or downplayed in light green in Table˜A.12 and Table˜A.13.
@@
SeViLA (Yu et al., 2024a) 32 61.3 61.5 75.6 63.6
VideoTree (Wang et al., 2024g) (56) 75.2 67.0 81.3 73.5
@@
SeViLA (Yu et al., 2024a) 32 - - - 60.9
VideoTree (Wang et al., 2024g) (56) - - - 66.9
@@
Table A.13: Extended results on IntentQA (Li et al., 2023). We compare LVNet against prior zero-shot models across different reasoning categories: Why?, How?, and B.A. (Belief/Action). LVNet achieves an overall accuracy of 71.7%, surpassing all models while using only 12 captions per video. It reaches 75.0% in the Why? category, 74.4% in the How? category, and 62.1% in the B.A. category. Compared to VideoTree, which processes 56 captions and achieves 66.9% accuracy, LVNet outperforms it by +4.8% while using significantly fewer captions. Additionally, LVNet demonstrates superior reasoning-based performance compared to LangRepo (90 captions, 59.1%) and LLoVi (90 captions, 64.0%). Models with video-caption pretraining or utilizing significantly more captions than 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency. Numbers in parentheses () indicate the maximum number of frames used.
Appendix A.3 Extended results on VideoMME
[…]
Qwen-VL (Bai et al., 2023) 7B/OS no no 4 N/A 37.8
Qwen-VL 7B/OS no no 1800 N/A OOM
LongVILA (Chen et al., 2024a) 8B/OS no no 128 N/A 38.8
LongVILA 8B/OS no no 256 N/A 39.7
LLaVA-OneVision (Li et al., 2024) 7B/OS no no 8 N/A 43.8
VideoChat-T (Zeng et al., 2024) 7B/OS no no 128 N/A 87.5 $\mathcal{O}(n)$ 41.9
Frame-Voyager (Yu et al., 2024b) 7B/OS no no 128 N/A 93.7 $\mathcal{O}{(n\textsuperscript{r})}$ 48.9
LLaVA-NexT-Video (Zhang et al., 2024a) 34B/OS no no 32 N/A 44.3
Frame-Voyager 34B/OS no no 128 N/A 93.7 $\mathcal{O}{(n\textsuperscript{r})}$ 51.2
InternVL2 (Chen et al., 2024c) 34B/OS no no 16 N/A 52.6
LVNet (DeepSeek-v3) 37B/OS yes yes 1800 24 $\mathcal{O}(n)$ 53.1
VideoAgent+GPT4o <1.8T/$P$ yes yes 24 $\mathcal{O}(m)$ 51.3
VideoTree+GPT4o <1.8T/$P$ yes yes 300 24 92.0 $\mathcal{O}{(n\textsuperscript{r})}$ 52.3
GPT4o direct <1.8T/$P$ yes yes 1800 1800 API Error
LVNet (GPT-4o) <1.8T/$P$ yes yes 1800 24 98.7 $\mathcal{O}(n)$ 53.9
LLaVA-Video (Zhang et al., 2024b) 72B/OS no no 64 N/A 61.5
Qwen2-VL (Wang et al., 2024c) 72B/OS no no 768 N/A 62.2
InternVL2.5 (Chen et al., 2024b) 72B/OS no no 64 N/A 62.6
Table A.14: Detailed Comparison on VideoMME (Fu et al., 2024) . This table expands on Table˜3 by including two-stage (TS), video-level training free (VT Free), number of input frames used, two-stage captions fed to LLM (TS Cap.), ratio of frames/captions provided to the LLM relative to input frames (FR), and frame selection complexity. “OS” denotes open-source, "OOM” stands for out-of-memory. $\mathcal{O}(m)$ denotes the complexity proportional to the number LLM calls to predict keyframe timelines. The bottom 3 results are from benchmark leaderboard; we are unable to replicate these on our compute resources. Our LVNet variants require no video-level training yet achieve competitive results on very long videos.
Table˜A.14 provides an extended comparison of both single-stage and two-stage approaches on the VideoMME long-video benchmark. In contrast to single-stage methods, which typically require costly video-level training on large datasets which has the risk of containing evaluation datasets. LVNet is video-level training free and thus retains broader generality. Critically, single-stage pipelines often become infeasible for long videos because they must process entire sequences—potentially exceeding 1,000 frames—through a heavy vision-language model (VLM) or LLM.
Keyframe-selection approaches mitigate this challenge by filtering a minimal set of relevant frames before any large-model processing. As shown in Table˜A.14, LVNet can handle up to 1,800 frames by reducing them to just a few dozen keyframes, thereby remaining both memory- and compute-efficient while still achieving strong accuracy. For Qwen-VL and GPT4o, we demonstrate an ablation with 1,800 frames to highlight how easily single-stage methods can run out of memory (OOM) under 4 NVIDIA RTX A5000 GPUs or encounter GPT API errors.
Moreover, our approach achieves strong performance with both open-source and proprietary LLM backbones, all without requiring any video-level training. LVNet+DeepSeek-V3 outperforms these single-stage keyframe selection models (VideoChat-T, Frame-Voyager) with equal or less frame selection complexity and outperforms single-stage VideoLLMs with a similarly sized open-source LLM. In a direct two-stage head-to-head setting that uses GPT-4o for both frame captioning and LLM reasoning and enforces the same 24-caption budget, LVNet further surpasses VideoAgent and VideoTree by +2.6% and +1.6%, respectively. Taken together, these results show LVNet’s effectiveness over both single-stage and two-stage video methods. The accuracy comes from each model’s recommended configuration and matched budgets for fair comparison.
Appendix A.4 Qualitative Analysis of Hierarchical Keyframe Selector
Figure A.6: Open-ended Responses from LVNet vs Uniform Sampling: The frames chosen by LVNet and the naive uniform sampling method are indicated with blue and red checkmarks, respectively. LVNet identifies both welding torches and measuring tapes, choosing the correct option, whereas uniform sampling only detects welding tools and selects the incorrect answer. The blue, red, and purple highlights correspond to the three main activities in the video—welding a handle, using a hammer, and using a measuring tape, respectively.
We compare open-ended responses of LVNet and the uniform sampling method in Figure˜A.6 to understand the effectiveness of the hierarchical keyframe selector in LVNet. The frames chosen by LVNet and the naive uniform sampling method are indicated by blue and red checkmarks in the images, respectively. LVNet selects frames at 5, 69, and 135 seconds by executing the hierarchical keyframe selector and generates captions based on those frames. When we feed the concatenated captions to the LLM to answer the given question: "Based on the video, what are the three main types of tools that C uses…" in an open-ended manner, the output identifies two main activities: welding torches and measuring tapes, among the three main activities described in Option 3 (welding handle, hammer, measuring tape), which is the correct answer, leading LVNet to choose the correct option.
@@ Appendix A.5 Algorithms in Detail
Our algorithms are presented in full detail in Algorithm˜1, Algorithm˜2, and Algorithm˜3. TSC in Algorithm˜1 extracts per-frame visual features using ResNet-18, followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the $n$ sets, we uniformly sample $\leq\tau$ frames, obtaining a total of $T_{a}\leq\tau\times n$ frames. For example, LVNet sets $\psi=5,\lambda=12,\tau=18$, resulting in approximately $n\sim 25$ and $T_{a}\sim 390$ on the EgoSchema dataset. CKD in Algorithm˜2 selects top $L$ frames based on similarity/confidence scores, which are calculated using cosine similarity between frames and keywords with CLIP-B/16. LVNet employs $L=32,len(K)\leq 25$ on the EgoSchema dataset. FKD in Algorithm˜3 sorts frames and their corresponding keywords by confidence scores, and reorder the $K$ frames with the lowest scores temporally. It groups frames sequentially into visual templates, each consisting of $N$ frames. From each template, the $M$ frames and keywords most relevant among the $N$ pairs are selected using GPT-4o. We set $L=32,K=16,N=8,M=3$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
1323:  $\delta$ $\leftarrow$ $\gamma~th$ value in $\textbf{List}_{index}$
@@
[…]
@@
[…]
@@
21:Require: keyword set $\mathbf{K}$, image set $\mathbf{I}$, similarity score list $\mathbf{S}$, total length $L$, number of low similarity indices $K$, number of frames per visual template NN, number of keyframes selected per visual template $M$, function to sort by similarity $\textbf{sort}(\mathbf{S})$, function to order indices temporally $\textbf{temporal\_order}()$
3
@@
[…]
@@
[…]
@@
1111:return $\mathbf{I}_{selected}$
@@ Appendix A.6 Prompting: Fine Keyframe Detector
We prompt the VLM to select frames that are most compatible with the list of given keywords. Each template image contains 8 images, and their order is described in language (e.g. top left to right, bottom left to right) and the VLM outputs the selected images according to our prompting as described in Figure˜A.7.
@@
Appendix A.7 Integration to Existing Methods
Our LVNet has been successfully integrated with other works for evaluation on long video benchmarks. For example, in Ranasinghe et al. (2024) LVNet is integrated with their proposed MVU to gain further performance boosts on the EgoSchema and NextQA benchmarks.
@@ Appendix A.8 Comparison with Other Keyframe Selection Methods
Figure˜5 visualizes the differences of the keyframe selection mechanism bewtween LVNet and VideoAgent. On the left, LVNet begins with uniformly sampled frames and filters them through multiple stages, resulting in a non-uniform distribution of frames over time. First, the temporal scene clustering (TSC) selects some frames that represent temporally distinct activities. Next, the coarse keyframe detector (CKD) targets frames most relevant to the question. Finally, the fine keyframe detector (FKD) further refines this selection to ensure the keyframes accurately capture the activity in question. As a result, LVNet produces 12 frames, with 8 of them (67%) directly depicting "usage of phones," which is the correct answer and leads the model to select the right option. On the right, VideoAgent also starts with the uniform frames but relies on a LLM to request additional frames. Since the initial frames do not capture enough relevant content, the LLM again selects frames uniformly, adding more irrelevant samples that lack the crucial information about "usage of phones." As a result, VideoAgent ultimately selects the wrong option.